# On Safety and Liveness Filtering Using Hamilton-Jacobi Reachability Analysis

### v1

@@ IV-C Smooth Least Restrictive Filter
While the LR filter (15) blends performance and liveness, it still switches abruptly to $\pi^{*}_{\text{live}}$ at the boundary of $\mathcal{G}_{\text{live}}$, leading to a high energy, bang-bang behavior near the boundary of $\mathcal{G}_{\text{live}}$. To overcome this challenge, we propose an alternative liveness filter that uses the full control authority available to the system as the projection set, rather than just $\{\pi^{*}_{\text{live}}\}$:
$\begin{split}\pi(x,t)=\arg\min_{u\in\mathcal{U}}&\quad||u-\pi_{\text{nom}}(x,t)||_{2}^{2}\\ s.t.&\quad u\in\mathcal{U}_{\text{live}}(x,t)\end{split}$ (17)
i.e., $\widetilde{\mathcal{U}}(x,t)=\mathcal{U}_{\text{live}}(x,t)$ for all $(x,t)$. Intuitively, the above optimization problem computes a control input at each time instant that is “closest” to the nominal controller, yet is within the set of live controls. Furthermore, since $\mathcal{U}_{live}=\mathcal{U}$ whenever $V(x,t)<0$, we can simplify the filter in (17) as:
$\pi(x,t)=\begin{cases}\pi_{\text{nom}}(x,t)&V(x,t)<0\\ \pi^{+}(x,t)&V(x)=0\end{cases}$ (18)
where $\pi^{+}(x,t)$ is obtained by solving the following optimization problem:
$\begin{split}\min_{u\in\mathcal{U}}&\quad||u-\pi_{\text{nom}}(x,t)||_{2}^{2}\\ s.t.&\quad D_{t}V(x,t)+\nabla V(x,t)\cdot f(x,u,d^{*})=0,\end{split}$ (19)
where $d^{*}$ is the optimal disturbance given by (10). We refer to the controller in (18) as the smooth least restrictive (smooth LR) filter. Similar to the LR filter (15), the smooth LR filter (18) follows $\pi_{\text{nom}}$ when the system state is in the interior of $\mathcal{G}_{\text{live}}(t)$. However, when the system state is at the boundary of $\mathcal{G}_{\text{live}}(t)$, rather than using $\pi^{*}_{\text{live}}(x,t)$, it chooses a live control that is closest to the nominal control by solving a Quadratic Program (QP), thus allowing for a smoother control profile.
@@
Running Example (Liveness). The trajectory and control profile corresponding to the smooth LR filter are shown in orange in Fig. 3 and 4. We observe that the smooth LR filter effectively “smoothes” out the bang-bang behaviors that the LR filter displayed; control profile now follows a gradual curve dictated by the QP formulation. This results in lower control energy (orange bar in Fig. 6) and less jittery behavior of the resulting system when compared to the original LR controller.
@@ IV-D Smooth Blending of Performance and Liveness
$\begin{split}\widetilde{\mathcal{U}}(x,t)=&\{u\in\mathcal{U}:\\ &D_{t}V(x,t)+\nabla V(x,t)\cdot f(x,u,d^{*})\leq-\gamma V(x,t)\}\end{split}$ (20)
The left hand side of the constraint in (20) is the total derivative of $V(x,t)$ evaluated at state $x$ and time $t$. Thus, intuitively, the above constraint limits the rate at which the value function can increase, i.e., approach the boundary of $\mathcal{G}_{\text{live}}(t)$, even when the system is farther from the boundary of $\mathcal{G}_{\text{live}}(t)$. When the system is at the boundary of the $\mathcal{G}_{\text{live}}(t)$, the RHS is zero and the controller must output a control that would keep the system at the same value. Specifically comparing with the LR and smooth LR filters in this section, the projection set in (20) blends $\pi_{\text{nom}}$ with $\pi^{*}_{\text{live}}$ in such a way that the system can take gradually more stringent corrective actions as the system becomes more at risk of breaching the liveness, resulting in less jerky controls.
@@
[…]
Intuitively, Lemma 3 states that (a) using the projection set in (20) always results in a feasible filtering problem, and (b) liveness is always guaranteed under the resultant control input, regardless of the value of $\gamma$. An alternative interpretation of Lemma 3 is that, as long as the value function $V(x,t)$ is differentiable, it acts as a robust CBF for any value of $\gamma$ and can be used to synthesize CBF-like controllers. Even though a number of works have synthesized CBFs using the reachability analysis (e.g., [9]), these methods need to explicitly select a value of $\gamma$ during the value function computation, making it challenging to tune $\gamma$ online. In contrast, Lemma 3 indicates that $\gamma$ can in fact be chosen online, independent of the value function computation.
@@
Effect of $\gamma$. To illustrate the effect of $\gamma$, we plot the overall (normalized) trajectory cost for different values of $\gamma$ in Fig. 5. As evident from the figure, the choice of $\gamma$ can significantly affect the performance of the smooth blending filter. Specifically, as $\gamma$ approaches 00, the system behaves more and more like $\pi^{*}_{\text{live}}$, keeping stringent liveness characteristics. This liveness, however, comes at the cost of a drop in performance (the controller cost is very high for small $\gamma$). On the other hand, as we increase $\gamma$ (till around $\gamma=1$), we see a proclivity towards using a controller That is close to the nominal controller while encouraging liveness. This results in a significant decrease in the trajectory cost. However, beyond a certain $\gamma$ ($\gamma=1$ in this case), the cost is again seen to rise slightly. This is because, for such a high $\gamma$, the controller is allowed to behave more and more like the smooth LR filter. The blending filter chooses the nominal controller as long as the system remains live; however, this choice leads the system closer to the BRT boundary, and the controller has to revert to a high-cost control in order to keep the system live, ultimately raising the overall controller cost. Hence, for high $\gamma$, the smooth-blending controller have a similar cost as the smooth LR controller.
@@ IV-E Comparison of Different Liveness Filters
    controller computation time per step, measuring the filter latency;
@@
    control energy, measuring how much control authority is used;
@@
    jerk energy in control, measuring the jerk in the control profile.
@@
Figure 6: Comparison of different liveness filters across performance, computation time, and smoothness of the control profile. MPC is the nominal controller. LR and SLR stand for the least-restrictive and the smooth least-restrictive filters. SB stands for smooth blending. We use the nominal controller as the reference normalizing factor. The nominal controller is normalized to 1 in the distance, cost, and energy computations and 0 in the time computations.
@@
First of all, we note that despite a poor nominal controller, all liveness filters steer the system very close to the target set (minimum distance to the target set is close to zero). However, the LR filters have a small non-zero distance to the target set. This is because in our simulation we use a small but non-zero timestep, which causes a small delay in the switching from the nominal to the liveness ensuring controller at the BRT boundary, resulting in the system state being outside of the BRT momentarily. We discuss this aspect further, as well as potential remedies in Sec. VII-A. Interestingly, we note that even though for high $\gamma$, we expect the smooth blending filter to behave similar to the smooth LR filter, we do not observe a similar loss of liveness in the smooth LR filter. This is because even for a high $\gamma$, the smooth blending filter applies the liveness controller before the system reaches the BRT boundary.
@@
In terms of performance, we note that $\pi^{*}_{\text{live}}$ results in a very high trajectory cost, which in our running example is dominated by a high control energy cost due to its bang-bang nature. In contrast, liveness filters balance liveness with performance, resulting in a lower trajectory cost. In this case, we note that smoother the control profile, the lower the trajectory cost, resulting in a better performance by smooth blending filter followed by smooth LR filter followed by LR filter. We further note that for $\gamma=50$, the performance of the smooth blending and smooth LR filters is very close, as expected. However, as $\gamma$ decreases, the smooth blending filter prioritizes liveness over performance, resulting in an increased trajectory cost. We see a very similar pattern in the control energy plots.
@@
Next, we compare the computation time of different controllers, normalized by the mean computation time for the $\pi_{\text{nom}}$. Here, $\pi^{*}_{\text{live}}$ emerges as the controller of choice, as it only requires computing $\pi_{\text{live}}^{*}$, which in typically is very fast, especially when the value function is represented as a neural network. In fact, in this case, computing the nominal controller is more expensive than computing $\pi_{\text{live}}^{*}$, which is why the mean computation time for the default controller is negative. For the same reason, the LR filter also leads to a negative computation cost, as it resorts to $\pi_{\text{live}}^{*}$ when the system state is at the BRT boundary. Compared to the aforementioned controllers, the smooth LR and smooth blending filters leads to a substantially higher computation time, as they require solving an optimization problem to compute the control input which constitutes most of their computation time. Moreover, the smooth LR filter has a much lower computation cost among the two because it only solves an optimization problem when the system state is at the BRT boundary. In general, the better the nominal controller is at maintaining liveness, the smaller the computation time for the smooth LR filter.
@@
Overall recommendation. Upon comparing different liveness filters, we recommend to use the smooth blending controller with a high value of $\gamma$, as it nicely tradeoff performance and liveness, while maintaining a low jerk. At the same time, using a very high $\gamma$ can spare the user from tuning of $\gamma$, which can significantly affect the performance of the system (Fig. 5). A close second choice is the smooth least restrictive filter, which also share the above advantages. In addition, it has a lower computation latency, making it particularly suitable for real-world robotic systems. One disadvantage of the smooth LR filter is that it can result in a loss of liveness if not implemented with a small enough timestep, which is a key reason behind our overall pick.
@@ V Safety Filtering Using Reachability Analysis
We now turn our attention to safety filtering. Similar to liveness filtering, we will first characterize the set of all safe control inputs for the system, and then use it to design various safety filters. As noted in Remark 1, we will use the converged value function for designing safety filters. Nevertheless, the proposed filters can easily be extended to incorporate time, similar to liveness filters.
@@ V-A Characterizing the set of safe controls
To ensure safety, we would like the system to stay out of the $\mathcal{G}_{\text{unsafe}}$ at all times. Thus, intuitively, the set of safe control $\mathcal{U}_{\text{safe}}(x)$ at a state $x$ are the control inputs that instantaneously keep the system state outside $\mathcal{G}_{\text{unsafe}}$, thereby ensuring recursive safety. Mathematically,
$\mathcal{U}_{\text{safe}}(x)=\left\{\begin{array}[]{l}\mathcal{U}\quad\text{ if }V(x)>0\\ \\ \{u:\nabla V(x)\cdot f(x,u,d^{*})=0\}\quad\text{ if }V(x)=0\\ \\ \emptyset\quad\text{ if }V(x)<0\end{array}\right.$ (21)
@@
$\begin{split}\pi(x,t)=\arg\min_{u}&\quad h(u,\pi_{\text{nom}}(x,t))\\ s.t.~{}&u\in\widetilde{\mathcal{U}}(x,t),\text{~{}with~{}}\widetilde{\mathcal{U}}(x,t)\subseteq\mathcal{U}_{\text{safe}}(x)\end{split}$ (22)

### v2

@@ IV-C Smooth Least Restrictive Filter
While the LR filter (14) accounts for performance and liveness, it still switches to a high energy, bang-bang policy $\pi^{*}_{\text{live}}$ at the boundary of $\mathcal{G}_{\text{live}}$. To overcome this challenge, we propose an alternative liveness filter that uses the full control authority available to the system as the projection set, rather than just $\{\pi^{*}_{\text{live}}\}$:
$\begin{split}\pi(\mathbf{x},t)=\arg\min_{u\in\mathcal{U}}&\quad||u-\pi_{\text{nom}}(\mathbf{x},t)||_{2}^{2}\\ s.t.&\quad u\in\mathcal{U}_{\text{live}}(\mathbf{x},t)\end{split}$ (16)
i.e., $\widetilde{\mathcal{U}}(\mathbf{x},t)=\mathcal{U}_{\text{live}}(\mathbf{x},t)$ for all $(\mathbf{x},t)$. Intuitively, the above optimization problem computes a control input at each time instant that is “closest” to the nominal controller, yet is within the set of live controls. Furthermore, since $\mathcal{U}_{live}=\mathcal{U}$ whenever $V(\mathbf{x},t)<0$, we can simplify the filter in (16) as:
$\pi(\mathbf{x},t)=\begin{cases}\pi_{\text{nom}}(\mathbf{x},t)&V(\mathbf{x},t)<0\\ \pi^{+}(\mathbf{x},t)&V(\mathbf{x},t)=0\end{cases}$ (17)
where $\pi^{+}(\mathbf{x},t)$ is obtained by solving the following optimization problem:
$\begin{split}\min_{u\in\mathcal{U}}&\quad||u-\pi_{\text{nom}}(\mathbf{x},t)||_{2}^{2}\\ s.t.&\quad D_{t}V(\mathbf{x},t)+\nabla V(\mathbf{x},t)\cdot f(\mathbf{x},u,d^{*})=0\end{split}$ (18)
where $d^{*}$ is the optimal disturbance given by (10). We refer to the controller in (17) as the smooth least restrictive (smooth LR) filter. Similar to the LR filter (14), the smooth LR filter (17) follows $\pi_{\text{nom}}$ when the system state is in the interior of $\mathcal{G}_{\text{live}}(t)$. However, when the system state is at the boundary of $\mathcal{G}_{\text{live}}(t)$, rather than using $\pi^{*}_{\text{live}}(\mathbf{x},t)$, it chooses a live control that is closest to the nominal control by solving a Quadratic Program (QP), thus allowing for a smoother control profile.
@@
Running Example (Liveness). The trajectory and control profile corresponding to the smooth LR filter are shown in orange in Fig. 2 and 3. We observe that the smooth LR filter effectively “smoothes” out the bang-bang behaviors that the LR filter displayed; control profile now follows a gradual curve dictated by the QP formulation. This results in lower control energy and less jittery behavior of the resulting system when compared to the original LR controller.
@@ IV-D Smooth Blending of Performance and Liveness
$\begin{split}\widetilde{\mathcal{U}}(\mathbf{x},t)=&\{u\in\mathcal{U}:\\ &D_{t}V(\mathbf{x},t)+\nabla V(\mathbf{x},t)\cdot f(\mathbf{x},u,d^{*})\leq-\gamma V(\mathbf{x},t)\}\end{split}$ (19)
where $\gamma$ is a tunable parameter that determines how quickly a control is permitted to drive the systems towards the boundary of $\mathcal{G}_{\text{live}}(t)$. The left hand side of the constraint in (19) is the total derivative of $V(\mathbf{x},t)$ evaluated at state $\mathbf{x}$ and time $t$. Thus, intuitively, the above constraint limits the rate at which the value function can increase, i.e., how quickly is the system allowed to approach the boundary of $\mathcal{G}_{\text{live}}(t)$, at any given state $\mathbf{x}$. When the system is on the boundary of $\mathcal{G}_{\text{live}}(t)$, we have $-\gamma V(\mathbf{x},t)=0$, and the controller must output a control that would drive the system to a state that has same value. Specifically comparing with the LR and smooth LR filters in this section, the projection set in (19) blends $\pi_{\text{nom}}$ with $\pi^{*}_{\text{live}}$ in such a way that the system can take gradually more stringent corrective actions as the system becomes more at risk of breaching the liveness, resulting in less jerky controls.
@@
Let $\mathbf{x}$ be any state such that $V(\mathbf{x},t)\leq 0$ and $\mathbf{x}\not\in\mathcal{L}$. Define $\widetilde{\mathcal{U}}(\mathbf{x},t)=\{u\in\mathcal{U}:D_{t}V(\mathbf{x},t)+\nabla V(\mathbf{x},t)\cdot f(\mathbf{x},u,d^{*})\leq-\gamma V(\mathbf{x},t)\}$. Then (a) $\widetilde{\mathcal{U}}(\mathbf{x},t)$ is non-empty, and (b) $\widetilde{\mathcal{U}}(\mathbf{x},t)\subseteq\mathcal{U}_{\text{live}}(\mathbf{x},t)$ for all $\gamma\geq 0$.
Intuitively, Lemma 3 states that (a) using the projection set in (19) always results in a feasible filtering problem, and (b) liveness is always guaranteed under the resultant control input, regardless of the value of $\gamma$. An alternative interpretation of Lemma 3 is that, as long as the value function $V(\mathbf{x},t)$ is differentiable, it acts as a robust CBF for any value of $\gamma$ and can be used to synthesize CBF-like controllers. Even though a number of works have synthesized CBFs using the reachability analysis (e.g., [10]), these methods need to explicitly select a value of $\gamma$ during the value function computation, making it challenging to tune $\gamma$ online. In contrast, Lemma 3 indicates that $\gamma$ can in fact be chosen online, independent of the value function computation.
@@
Effect of $\gamma$. To illustrate the effect of $\gamma$, we plot the overall trajectory cost (given by (3)), normalized by the cost of the default controller for different values of $\gamma$ in Fig. 4. As evident from the figure, the choice of $\gamma$ can significantly affect the performance of the smooth blending filter. Specifically, as $\gamma$ approaches 00, the system behaves more and more like $\pi^{*}_{\text{live}}$, keeping stringent liveness characteristics. This liveness, however, comes at the cost of a drop in performance (the controller cost is very high for small $\gamma$). On the other hand, as we increase $\gamma$ (till around $\gamma=1$), we see a proclivity towards using a controller that is close to the nominal controller while encouraging liveness. This results in a significant decrease in the trajectory cost. However, beyond a certain $\gamma$ ($\gamma=1$ in this case), the cost is again seen to rise slightly. This is because, for such a high $\gamma$, the controller is allowed to behave more and more like the smooth LR filter. The blending filter chooses the nominal controller as long as the system remains live; however, this choice leads the system closer to the BRT boundary, and the controller has to revert to a high-cost control in order to keep the system live, ultimately raising the overall controller cost. Hence, for high $\gamma$, the smooth-blending controller has a similar cost as the smooth LR controller.
@@ IV-E Comparison of Different Liveness Filters
    computation time per step, measuring the filter latency;
@@
    control energy, measuring control authority used;
@@
    jerk energy, measuring the jerk in the control profile.
@@
First, we note that despite a poor nominal controller, all liveness filters steer the system very close to the target set (minimum distance to the target close to zero). However, most filters have a small non-zero distance to the target set. This is because in our simulation we use a small but non-zero timestep, which introduces a small delay in the application of the filtered policy, resulting in the system state being outside of the BRT momentarily. We discuss this aspect further, as well as potential remedies in Sec. VII-A. The exception is the smooth blending filter with small $\gamma$, as it very closely behaves like the default controller sacrificing performance almost entirely to reach the target.
Figure 5: Comparison of different liveness filters across performance, computation time, and smoothness of the control profile. MPC is the nominal controller. LR and SLR stand for the least-restrictive and the smooth least-restrictive filters. SB stands for smooth blending.
@@
In terms of performance, we note that $\pi^{*}_{\text{live}}$ results in a very high trajectory cost, mostly due a high control energy cost expected of its bang-bang nature. In contrast, liveness filters balance liveness with performance, resulting in a lower trajectory cost. In this case, we observe that smoother control profiles typically equate to lower trajectory costs, resulting in a better performance by smooth blending filter followed by smooth LR filter followed by LR filter. We further note that for $\gamma=50$, the performance of the smooth blending and smooth LR filters is very close, as expected. However, as $\gamma$ decreases, the smooth blending filter prioritizes liveness over performance, resulting in an increased trajectory cost. We see a very similar pattern in the control energy plots.
@@
Next, we compare the computation time of different controllers. Here, the default controller emerges as the fastest option, as it only requires computing $\pi_{\text{live}}^{*}$, which typically is very fast, especially when the value function is represented as a neural network. In fact, in this case, computing the nominal controller is more expensive than computing $\pi_{\text{live}}^{*}$. For the same reason, the LR filter is faster than the MPC method, as it resorts to $\pi_{\text{live}}^{*}$ when the system state is at the BRT boundary. Compared to the aforementioned controllers, the smooth LR and smooth blending filters leads to a higher computation time, as they require solving an optimization problem additionally to computing the nominal MPC controller. Moreover, the smooth LR filter has a lower computation cost among the two because it only solves an optimization problem when the system state is at the BRT boundary. In general, the better the nominal controller is at maintaining liveness, the smaller the computation time for the smooth LR filter.
@@
Overall recommendation. Upon comparing different liveness filters, we recommend to use the smooth blending controller with a high value of $\gamma$, as it nicely trades off performance and liveness trough the tuning of $\gamma$, while maintaining a low jerk. At the same time, it allows to pick a large $\gamma$ to spare the user from tuning while keeping a considerable performance increase as seen in Fig. 4. A close second choice is the smooth least restrictive filter, which also shares the above advantages. In addition, it has a lower computation latency, making it particularly suitable for real-world robotic systems.
@@ V Safety Filtering Using Reachability Analysis
We now turn our attention to safety filtering. One important difference to note is that the value function often converges for the safety problem, i.e. $\mathcal{G}_{\text{unsafe}}$ stops growing after some amount of time, beyond which the system has enough time to avoid the target set despite the worst case disturbance. Consequently, we can use the converged value function $V(\mathbf{x})$, which is no longer a function of time $t$, to synthesize safety controllers, and the resulting controllers are time-invariant and can be expressed as:
$\pi_{\text{safe}}^{*}(\mathbf{x})=\arg\max_{u\in\mathcal{U}}\min_{d\in\mathcal{D}}\nabla V(\mathbf{x})\cdot f(\mathbf{x},u,d)$ (20)
Nevertheless, the proposed filters can easily be extended to incorporate time, similar to liveness filters. To demonstrate its effectiveness, we introduce the following example case.
Running Example (Safety). Consider a 4D system model of a blimp with state $\mathbf{x}=(x,y,z,\theta)$, with control over the rate of change of altitude and yaw angle, as described by the following dynamics:
$\dot{x}=v\cos(\theta)+{d}_{1}$ (21)
$\dot{y}=v\sin(\theta)+{d}_{2}$
$\dot{z}=u_{1}$
$\dot{\theta}=u_{2}$
where $x,y$ and $z$ denote the $X,Y$ and $Z$ positions of the center of mass of the system, respectively, and $v$ is the speed of the system on the $XY$ plane. The control inputs are the vertical velocity in the $Z$ direction and the rate of change of heading. In addition, there are velocity disturbances in the $X$ and $Y$ directions. We consider a $10m$ x $10m$ x $5m$ position state space (Fig. 6) where the objective is to reach a target area of radius $R=0.5m$ located at $(x,y,z)=(7,7,3)m$ (pink sphere) while avoiding collision with the obstacle set $\mathcal{L}$ that corresponds to a collection of spheres of various sizes spread across the position state space (gray spheres). In this case, the implicit obstacle function is defined as the signed distance to the closest gray sphere, $l(\mathbf{x})=\min_{i}{d(\mathbf{x},S_{i})}$ where $d(\mathbf{x},S_{i})=\sqrt{(x_{i}-x)^{2}+(y_{i}-y)^{2}+(z_{i}-z)^{2}}-R_{i}$ is the distance from state $\mathbf{x}$ to the sphere $S_{i}$ of radius $R_{i}$ centered at $(x_{i},y_{i},z_{i})$. In (21) $v=2m/s$ is the velocity in the $XY$ plane, the control and disturbance bounds are given as $u_{1}\in[-1,1]m/s$, $u_{2}\in[-\pi,\pi]\text{rad}/s$ and $||\begin{bmatrix}d_{1}&d_{2}\end{bmatrix}^{\top}||_{2}\leq 0.5m/s$.
For this example, to illustrate that the proposed method is agnostic to how the value function is obtained, we leverage the Level Set Toolbox [21] and its reachability helper library HelperOC [9] to compute a numerical approximation of the BRT. Level-Set methods solve the HJB-VI numerically over a uniformly discretized state space grid. We use a 4D grid of $81\times 81\times 41\times 21$ points for our computations. The solution of the HJB-VI provides us with the value function, $V(\mathbf{x},\tau)$, which can be used to extract the $\mathcal{G}_{\text{unsafe}}$ similar to (8). In this case, the value function converges after 0.5 seconds; thus, we consider the time converged $\mathcal{G}_{\text{unsafe}}$. A 3D slice of $\mathcal{G}_{\text{unsafe}}$ for a fixed value of $\theta=0$ is presented in Fig. 6. The expansion of the BRT can be seen in the opposite direction, as states oriented this way will inevitably collide with obstacles if they start too close. By definition of the BRT, as long as the system starts outside the blue region, it is guaranteed to remain safe under the default safety controller $\pi_{\text{safe}}^{*}$ (20).
Figure 6: A $xyz$-slice (for $\theta=0$) of the BRT, $\mathcal{G}_{\text{unsafe}}$, for the safety example is shown in blue. The gray spheres inside the BRT correspond to the obstacles (the target set in this case) and the pink sphere represents the set of goal states. The system trajectory under the default safety controller is shown in dark blue and the corresponding controller profiles are shown in the inlets.
The system trajectory under $\pi_{\text{safe}}^{*}$ starting at a state outside the BRT and the corresponding control profiles are shown in Fig. 6. The system remains safe under the default safety controller as expected; however, the default safety controller steers the system as far away from the obstacles as possible, disregarding any other performance criterion the system might have (such as reaching the pink goal region). Moreover, the control profile is jittery, which is typically undesirable for real-world applications.
@@ V-A Characterizing the set of safe controls
Similar to liveness filtering, we will first characterize the set of all safe control inputs for the system, and then use it to design various safety filters. To ensure safety, we would like the system to stay out of the $\mathcal{G}_{\text{unsafe}}$ at all times. Thus, intuitively, the set of safe control $\mathcal{U}_{\text{safe}}(\mathbf{x})$ at a state $\mathbf{x}$ are the control inputs that instantaneously keep the system state outside $\mathcal{G}_{\text{unsafe}}$, thereby ensuring recursive safety. Mathematically,
$\mathcal{U}_{\text{safe}}(\mathbf{x})=\left\{\begin{array}[]{l}\mathcal{U}\quad\text{ if }V(\mathbf{x})>0\\ \\ \{u\in\mathcal{U}:\nabla V(\mathbf{x})\cdot f(\mathbf{x},u,d^{*})=0\}\\ \text{ if }V(\mathbf{x})=0\\ \\ \emptyset\quad\text{ if }V(\mathbf{x})<0\end{array}\right.$ (22)
@@
Figure 7: Trajectories for the nominal control, the proposed safety filters, and a CBF-based filter.
@@
$\begin{split}\pi(\mathbf{x},t)=\arg\min_{u}&\quad h(u,\pi_{\text{nom}}(\mathbf{x},t))\\ s.t.~{}&u\in\widetilde{\mathcal{U}}(\mathbf{x},t),\text{~{}with~{}}\widetilde{\mathcal{U}}(\mathbf{x},t)\subseteq\mathcal{U}_{\text{safe}}(\mathbf{x})\end{split}$ (23)